Tessa can ride her bike 24 miles in one hour less than it took Gary to walk 9 miles.Tessa rate is 9 miles per hour faster than Gary rate find Tessa rate?
Answer by mananth(16949)

Tessa can ride her bike 24 miles in one hour less than it took Gary to walk 9 miles.Tessa rate is 9 miles per hour faster than Gary rate find Tessa rate?
...
let gary's rate be x mph
tessa's rate = y
..
y =x+9.............1
..
9/x-24/y =1
plug y = x+9
9/x-24/(x+9)=1
(9(x+9)-24x)/x(x+9)=1
9x+81-24x=x^2+9x
x^2+24x-81=0
x^2+27x-3x-81=0
x(x+27)-3(x+27)=0
(x+27)(x-3)=0
x= 3
gary's rate = 3mph.
y= x+9
plug value of x
y =3+9 = 12 mph. Tessa's rate
...
m.ananth@hotmail.ca